# Are We Overlooking the Dimensions? Learning Latent Hierarchical Channel Structure for High-Dimensional Time Series Forecasting

### v1

@@ 4 U-Cast Framework
Our theoretical and empirical analyses in Section 3 highlight the need for novel CD models that can efficiently handle high-dimensional time series. Traditional models often struggle with scalability and overlook the structured nature of such data, especially hierarchical correlations among variable groups. To address this, we propose U-Cast (see Figure 2), an efficient model that captures channel correlations via learning latent hierarchical structures. We also introduce a full-rank regularization term to encourage disentanglement and improve the learning of structured representations.
@@ 4.1 Model Architecture
Hierarchical Latent Query Network. Self-attention is well suited to modelling channel-wise dependencies because every variable can attend to every other variable without assuming locality. However, applying full attention to $C$ channels incurs quadratic cost, which is prohibitive when $C$ is large. Hierarchical latent query network addresses this by introducing a set of latent queries that serve as an information bottleneck. At layer $\ell\in\{1,\dots,L\}$ the number of latent queries is $C_{\ell}=\lfloor C/r^{\,\ell}\rfloor$, where $r>1$ is a pre-defined reduction ratio. The queries are shared across samples: $\mathbf{Q}_{\ell}\in\mathbb{R}^{C_{\ell}\times d}$. The Latent Query Attention at layer $\ell$ is
@@
followed by layer normalisation. Successive layers hierarchical latent query network builds a hierarchy in which higher-level queries summarise wider channel groups.
Temporal Alignment. The deepest latent representation $\mathbf{H}^{(L)}\in\mathbb{R}^{C_{L}\times d}$ stores $C_{L}$ latent channel tokens whose $d$-dimensional feature vectors encode the temporal dynamics extracted from the look-back window. To keep these temporal features coherent when the model switches from down-sampling to up-sampling, we apply a shared linear predictor $f_{\text{pred}}$ along the temporal dimension:
@@
Hierarchical Upsampling Network. Forecasting demands channel-wise outputs of size $C$, so the hierarchy is traversed in reverse to restore resolution. At layer $\ell=L,\dots,1$ the Up-Latent Query Attention uses the representation from the encoder as queries and the current decoder representation as keys and values:
@@
Output Projection. After upsampling $\mathbf{U}^{(0)}\in\mathbb{R}^{C\times d}$ has the same channel dimension as the encoder output $\mathbf{H}^{(0)}$. A residual link followed by a projection yields the horizon-length prediction:
@@ 4.2 Optimization Objective
The proof is given in Appendix G. Theorem 3 states that enforcing full rank on $\mathbf{H}^{(\ell)}$ is sufficient to remove linear redundancy among channels and reveal a clear hierarchical latent channel structure. Define the row-covariance matrix $\mathbf{\Sigma}^{(\ell)}=\tfrac{1}{d}\,\mathbf{H}^{(\ell)}\mathbf{H}^{(\ell)\!\top}\in\mathbb{R}^{C^{\prime}\times C^{\prime}}.$ The log-determinant $\log\det\bigl{(}\mathbf{\Sigma}^{(\ell)}+\varepsilon I_{C^{\prime}}\bigr{)}$ is proportional to the generalised variance [^1] Product of the eigenvalues of $\mathbf{\Sigma}^{(\ell)}$.. Maximising this value keeps every eigenvalue bounded away from zero, so the channel vectors occupy a larger subspace and share less redundant information. We thus define the full-rank regularization loss as
@@
where $\mathcal{L}_{\text{mse}}$ is the supervised mean-squared forecasting error, $\mathcal{L}_{\text{cov}}$ is the layer-averaged covariance penalty, and $\alpha$ controls the strength of this regulariser.
@@ 5 Time-HD: High-Dimensional Time Series Forecasting Benchmark
Table 2: Time-HD descriptions. The calculation of Correlation is illustrated in Appendix L.
@@
The goal of Time-HD is to support and advance research in time series forecasting, which is a rapidly growing and increasingly important area. Details on datasets generation, collection, and cleaning are provided in Appendix K. As shown in Table 2 and Figure 1, Time-HD has several key characteristics:
@@
Channel Correlation Statistic. We follow TFB qiu2024tfb to compute channel correlation. The results are shown in the “Correlation” column of Table 2. All datasets exhibit high average channel correlation, confirming the presence of strong inter-variable dependencies in these high-dimensional time series. The details of correlation calculation is illustrated in Appendix L.
Table 3: Forecasting results on various datasets. The prediction length of each dataset is described in Table 2. A lower MSE or MAE indicates better performance. The best performance is highlighted in red, and the second-best is underlined.
@@ 6 Experiment Results and Framework Analysis
    Channel-Independent. We include three methods that employ a shared model backbone across all channels: DLinear dlinear , PAttn pattn , and PatchTST patchtst . These methods use linear projection, attention mechanisms, and the Transformer architecture, respectively, to model temporal dependencies.
@@
    Channel-Dependent. We adopt five methods that incorporate dedicated modules to capture inter-channel dependencies. TimesNet timesnet and TSMixer tsmixer are position-wise methods that capture channel correlations at each time step by embedding. iTransformer itransformer is a token-wise method that treats the entire time series of each channel as a token and models inter-channel dependencies using a Transformer. Additionally, we consider two clustering-based methods: CCM ccm and DUET duet , which explicitly group highly correlated channels into clusters and model dependencies within each cluster.
@@
Figure 3: MSE vs. average training time per batch on Wind dataset ($C=3{,}850$); marker size reflects GPU memory usage.
@@
Benchmark Results. The experimental setup is detailed in Appendix M. The results are shown in Table 3. We summarize the following observations from our benchmark results: First, although CI methods cannot explicitly model channel correlations, they are designed to learn shared temporal patterns across channels. This can help reduce the risk of overfitting, enabling CI methods to achieve competitive performance. Among them, PAttn generally achieves better results, indicating the effectiveness of attention mechanisms for temporal modeling. Second, among CD methods, iTransformer significantly outperforms others, suggesting that treating each channel as a token and using attention to model inter-channel dependencies is more effective than position-wise or cluster-based approaches in high-dimensional settings. We argue that the token-wise design helps mitigate noise in the temporal dimension and fully utilizes the benefits of the attention mechanism. Third, CD methods do not show a consistent or significant performance advantage over CI methods. However, our theoretical analysis in Section 3 suggests that CD can reduce the MSE risk compared to CI in multivariate time series forecasting. We believe this is due to current CD baselines either failing to effectively exploit complex channel dependencies or being unable not scale to high-dimensional time series.
@@
Performance Gains of U-Cast. As shown in Table 3, U-Cast achieves top performance on the high-dimensional Time-HD benchmark, ranking first on 11 datasets for MSE and 10 for MAE. This underscores its strong ability to capture complex inter-channel dependencies by learning hierarchical latent channel structures compared to other CD models. In contrast, iTransformer, the representative CD model that explicitly models multivariate correlations via Transformer, underperforms on many datasets. This may be due to the extreme complexity of channel dependencies in a high-dimensional setting, where the lack of an explicitly hierarchical structure learning in its attention mechanism limits its effectiveness. By explicitly learning latent hierarchical structure by hierarchical latent query and full-rank regularization, U-Cast is better equipped to handle such complexity.
@@
Model Efficiency. Beyond its strong forecasting performance, another notable advantage of U-Cast is its lightweight architecture. Figure 3 shows U-Cast achieves the best trade-off between training speed, memory footprint, and performance. We also find that U-Cast consistently achieves favorable training efficiency across varying dimensionalities and as dimensionality increases, the efficiency advantage of U-Cast over other methods becomes more pronounced (see Appendix O). Additionally, we conduct computational complexity analysis, which indicates that U-Cast lowers both time and memory by a factor of $r$ compared with iTransformer, yet still retains the expressive power of attention through its latent-query hierarchy (see Appendix J).
@@
7 Conclusion and Future Opportunities
In this paper, we propose U-Cast to address the unique challenges of effectively modeling these latent hierarchical channel structures within High-Dimensional Time Series Forecasting (HDTSF). To facilitate research in HDTSF and provide a testbed for U-Cast and future innovative approaches, we curate the Time-HD benchmark. Through extensive experiments using our new benchmark, we identify and highlight key research directions and opportunities within high-dimensional time series forecasting: (1) scalable modeling for high-dimensional time series forecasting, (2) enhanced inter-channel correlation learning, and (3) standardized benchmarking protocols. It is our aspiration that U-Cast, Time-HD, and their accompanying resources will serve as a significant catalyst for innovation and progress in the field.
@@ Appendix A Comparison with Existing Works
   M3 [32]
   M4 [33]
   LTSF-Linear [53]
   TSlib [48]
   BasicTS [21]
   BasicTS+ [21]
   Monash [14]
   Libra [2]
   TFB [38] ✓ (only one) ✓ (only one) ✓ (only one)
@@
   Chronos [1]
   Moment [16]
   Moirai [47]
   TimesFM [11]
   Time-MoE [40] ✓ (only one) ✓ (only one) ✓ (only one)
@@
High-Dimensional Time Series Datasets. As shown in Table 5, most existing benchmarks do not include any high-dimensional datasets. The only exception is TFB [38], which incorporates only one dataset with 2000 channels (Wiki2000) for evaluation. Among foundation models, although all of them include high-dimensional time series datasets, these datasets are primarily used for pretraining rather than evaluation. Time-MoE [40] is the only one that evaluates on a high-dimensional dataset (Global Temp) with 1000 channels. Both TFB and Time-MoE attempt to explore the high-dimensional regime, but they each include only one such dataset. This limited coverage constrains their effectiveness as comprehensive benchmarks.
Multivariate Time Series Forecasting. Channel-independent (CI): The CI modeling strategy uses a shared model backbone for all channels and processes each channel independently during the forward pass [35, 53, 51, 44, 42, 10, 23]. This design generally results in lower model capacity but offers greater robustness. However, since CI models treat each channel independently, they are unable to capture inter-channel correlations explicitly.
Channel-dependent (CD): In contrast, the CD modeling strategy incorporates modules that are specifically designed to capture dependencies across channels. These methods tend to offer higher model capacity and can leverage cross-channel information more effectively. Based on how the inter-channel correlations are handled, CD approaches can be categorized as follows: (1) Position-wise [7, 48, 49, 58, 59, 54]: These methods first project the channel dimension of each time step into a hidden embedding space, resulting in one embedding per time step. This approach may ignore correlations across channels at different time steps and could introduce noise. After generating the embeddings, these methods mainly focus on modeling temporal dependencies. (2) Token-wise [25, 46]: These methods treat each channel as a token and input the sequence of tokens into a Transformer model. This helps reduce irrelevant noise for forecasting. Although multi-layer Transformers have the potential to capture complex hierarchical inter-channel structures, they are extremely time-consuming and do not scale well to high-dimensional time series. Moreover, the lack of an explicit mechanism for learning hierarchical channel structure may limit their effectiveness. (3) Cluster-wise [6, 39]: These approaches divide the multivariate time series into disjoint channel clusters using clustering algorithms. However, using a single-layer clustering step may overlook the hierarchical organization present in high-dimensional time series, which can lead to suboptimal performance.

### v2

@@ 4 U-Cast Framework
Our theoretical and empirical analyses in Section 3 highlight the need for novel CD models that can efficiently handle high-dimensional time series. Traditional models often struggle with scalability and overlook the structured nature of such data, especially the hierarchical correlations among variable groups that frequently occur in real-world datasets (see Appendix M). To address this, we propose U-Cast (see Figure 1), an efficient model that captures channel correlations via learning latent hierarchical structures. We also introduce a full-rank regularization term to encourage disentanglement and improve the learning of structured representations.
@@ 4.1 Model Architecture
Hierarchical Latent Query Network. Self‑attention is well suited to modelling channel‑wise dependencies because every variable can attend to every other variable without assuming locality. However, applying full attention to $C$ channels incurs quadratic cost, which is prohibitive when $C$ is large. Hierarchical latent query network addresses this by introducing a set of latent queries that serve as an information bottleneck. At layer $\ell\in\{1,\dots,L\}$ the number of latent queries is $C_{\ell}=\lfloor C/r^{\,\ell}\rfloor$, where $r>1$ is a pre‑defined reduction ratio. The queries, which are learnable, are shared across samples: $\mathbf{Q}_{\ell}\in\mathbb{R}^{C_{\ell}\times d}$. The Latent Query Attention at layer $\ell$ is
@@
followed by layer normalisation. Successive layers hierarchical latent query network builds a hierarchy in which higher‑level queries summarise wider channel groups.
Temporal Alignment. The deepest latent representation $\mathbf{H}^{(L)}\in\mathbb{R}^{C_{L}\times d}$ stores $C_{L}$ latent channel tokens whose $d$-dimensional feature vectors encode the temporal dynamics extracted from the look‑back window. To keep these temporal features coherent when the model switches from down‑sampling to up‑sampling, we apply a shared linear predictor $f_{\text{pred}}$ along the temporal dimension:
@@
Hierarchical Upsampling Network. Forecasting demands channel‑wise outputs of size $C$, so the hierarchy is traversed in reverse to restore resolution. At layer $\ell=L,\dots,1$ the Up‑Latent Query Attention uses the representation from the encoder as queries and the current decoder representation as keys and values:
@@
Output Projection. After upsampling $\mathbf{U}^{(0)}\in\mathbb{R}^{C\times d}$ has the same channel dimension as the encoder output $\mathbf{H}^{(0)}$. A residual link followed by a projection yields the horizon‑length prediction:
@@ 4.2 Optimization Objective
The proof is given in Appendix I. Theorem 3 states that enforcing full rank on $\mathbf{H}^{(\ell)}$ is sufficient to remove linear redundancy among channels and reveal a clear hierarchical latent channel structure. Define the row-covariance matrix $\mathbf{\Sigma}^{(\ell)}=\tfrac{1}{d}\,\mathbf{H}^{(\ell)}\mathbf{H}^{(\ell)\!\top}\in\mathbb{R}^{C^{\prime}\times C^{\prime}}.$ The log-determinant $\log\det\bigl(\mathbf{\Sigma}^{(\ell)}+\varepsilon I_{C^{\prime}}\bigr)$ is proportional to the generalised variance [^1] Product of the eigenvalues of $\mathbf{\Sigma}^{(\ell)}$.. Maximising this value keeps every eigenvalue bounded away from zero, so the channel vectors occupy a larger subspace and share less redundant information. We thus define the full-rank regularization loss as
@@
where $\mathcal{L}_{\text{mse}}$ is the supervised mean‑squared forecasting error, $\mathcal{L}_{\text{cov}}$ is the layer‑averaged covariance penalty, and $\alpha$ controls the strength of this regulariser.
@@ 5 Time-HD: High-Dimensional Time Series Forecasting Benchmark
Table 3: Time-HD descriptions. Details on dataset generation, collection, cleaning, and prediction length determination are provided in Appendix D.
@@
The goal of Time-HD is to support and advance research in TSF, which is a rapidly growing and increasingly important area. As shown in Table 3, Time-HD has several key characteristics:
@@
Channel Correlation Statistic. We compute channel correlation following TFB (Qiu et al., 2024a), with results reported in the correlation column of Table 3. All datasets exhibit strong channel dependencies, which are not intentionally curated but naturally arise in large-scale real-world systems. The high correlation reflects complex inter-channel dependencies and a shared global trend across channels. Further analysis on channel correlation of Time-HD is provided in Appendix E.
Table 4: Forecasting results on various datasets. Input length $T$ is tuned per model for best performance (see Appendix F), and prediction length $S$ is given in Table 3. Lower MSE/MAE indicates better performance; best results are in red and second-best are underlined.
@@ 6 Experiment Results and Framework Analysis
    Channel-Independent. We include three methods that employ a shared model backbone across all channels: DLinear (Zeng et al., 2023), PAttn (Tan et al., 2024), and PatchTST (Nie et al., 2023). These methods use linear projection, attention mechanisms, and the Transformer architecture, respectively, to model temporal dependencies.
@@
    Channel-Dependent. We adopt five methods that incorporate dedicated modules to capture inter-channel dependencies. TimesNet (Wu et al., 2022) and TSMixer (Chen et al., 2023) are position-wise methods that capture channel correlations at each time step by embedding. iTransformer (Liu et al., 2024a) is a token-wise method that treats the entire time series of each channel as a token and models inter-channel dependencies using a Transformer. Additionally, we consider two clustering-based methods: CCM (Chen et al., 2024) and DUET (Qiu et al., 2024b), which explicitly group highly correlated channels into clusters and model dependencies within each cluster.
@@
Figure 2: MSE vs. average training time per batch on Wind dataset (C=3,850C=3{,}850); marker size reflects GPU memory usage.
@@
Benchmark Results. The experimental setup is detailed in Appendix F. The results are shown in Table 4. We summarize the following observations from our benchmark results: First, although CI methods cannot explicitly model channel correlations, they are designed to learn shared temporal patterns across channels. This can help reduce the risk of overfitting, enabling CI methods to achieve competitive performance. Among them, PAttn generally achieves better results, indicating the effectiveness of attention mechanisms for temporal modeling. Second, among CD methods, iTransformer significantly outperforms others, suggesting that treating each channel as a token and using attention to model inter-channel dependencies is more effective than position-wise or cluster-based approaches in high-dimensional settings. We argue that the token-wise design helps mitigate noise in the temporal dimension and fully utilizes the benefits of the attention mechanism. Third, CD methods do not show a consistent or significant performance advantage over CI methods. However, our theoretical analysis in Section 3 suggests that CD can reduce the MSE risk compared to CI in multivariate time series forecasting. We believe this is due to current CD baselines either failing to effectively exploit complex channel dependencies or being unable not scale to high-dimensional time series.
@@
Performance Gains of U-Cast. As shown in Table 4, U-Cast achieves top performance on the high-dimensional Time-HD benchmark, ranking first on 12 datasets for MSE and 11 for MAE, and further achieves an average forecasting error reduction of 15% compared to iTransformer (t-test p-value=1.34*10-5). While iTransformer, the representative CD model that explicitly models multivariate correlations via Transformer, serves as a strong baseline, it underperforms on many datasets compared to U-Cast. This highlights the advantage of learning hierarchical latent channel structures for capturing complex inter-channel dependencies. The limitation of iTransformer may stem from the extreme complexity of channel dependencies in high-dimensional settings, where the absence of explicitly hierarchical structure learning in its attention mechanism constrains its effectiveness. In contrast, by incorporating hierarchical latent queries and full-rank regularization, U-Cast is better equipped to handle such complexity.
@@
Model Efficiency. Beyond its strong forecasting performance, another notable advantage of U-Cast is its lightweight architecture. Figure 2 shows U-Cast achieves the best trade-off between training speed, memory footprint, and performance. We also find that U-Cast consistently achieves favorable training efficiency across varying dimensionalities and as dimensionality increases, the efficiency advantage of U-Cast over other methods becomes more pronounced (see Appendix Q). Additionally, we conduct computational complexity analysis, which indicates that U-Cast lowers both time and memory by a factor of $r$ compared with iTransformer, yet still retains the expressive power of attention through its latent‑query hierarchy (see Appendix L).
@@
7 Conclusion
In this paper, we propose U-Cast to address the unique challenges of effectively modeling these latent hierarchical channel structures within High-Dimensional Time Series Forecasting (HDTSF). To facilitate research in HDTSF and provide a testbed for future innovative approaches, we curate the Time-HD benchmark. Through extensive experiments, we identify and highlight key research directions and opportunities in Appendix B. It is our aspiration that U-Cast, Time-HD, Time-HD-lib, and their accompanying resources will serve as a significant catalyst for innovation and progress in the time series community (see Appendix C for broader impact).
@@ Appendix A Comparison with Existing Works
   M3 (Makridakis & Hibon, 2000)
   M4 (Makridakis et al., 2018)
   LTSF-Linear (Zeng et al., 2023)
   TSlib (Wu et al., 2022)
   BasicTS (Liang et al., 2022)
   BasicTS+ (Liang et al., 2022)
   Monash (Godahewa et al., 2021)
   Libra (Bauer et al., 2021)
   TFB (Qiu et al., 2024a) ✓ (only one) ✓ (only one) ✓ (only one)
@@
   Chronos (Ansari et al., 2024)
   Moment (Goswami et al., 2024)
   Moirai (Woo et al., 2024)
   TimesFM (Das et al., 2024)
   Time-MoE (Shi et al., 2024) ✓ (only one) ✓ (only one) ✓ (only one)
@@
Figure 4: Time-HD provides diverse high-dimensional datasets.
High-Dimensional Time Series Datasets. As shown in Table 6, most existing benchmarks do not include any high-dimensional datasets. The only exception is TFB (Qiu et al., 2024a), which incorporates only one dataset with 2000 channels (Wiki2000) for evaluation. Among foundation models, although all of them include high-dimensional time series datasets, these datasets are primarily used for pretraining rather than evaluation. Time-MoE (Shi et al., 2024) is the only one that evaluates on a high-dimensional dataset (Global Temp) with 1000 channels. Both TFB and Time-MoE attempt to explore the high-dimensional regime, but they each include only one such dataset. This limited coverage constrains their effectiveness as comprehensive benchmarks.
Multivariate Time Series Forecasting. Channel-independent (CI): The CI modeling strategy uses a shared model backbone for all channels and processes each channel independently during the forward pass (Nie et al., 2023; Zeng et al., 2023; Xu et al., 2023; Wang et al., 2024a; Tan et al., 2024; Dai et al., 2024; Lin et al., 2024). This design generally results in lower model capacity but offers greater robustness. However, since CI models treat each channel independently, they are unable to capture inter-channel correlations explicitly.
Channel-dependent (CD): In contrast, the CD modeling strategy incorporates modules that are specifically designed to capture dependencies across channels. These methods tend to offer higher model capacity and can leverage cross-channel information more effectively. Based on how the inter-channel correlations are handled, CD approaches can be categorized as follows: (1) Position-wise (Chen et al., 2023; Wu et al., 2022; 2021; Zhou et al., 2021; 2022; Zhang & Yan, 2023): These methods first project the channel dimension of each time step into a hidden embedding space, resulting in one embedding per time step. This approach may ignore correlations across channels at different time steps and could introduce noise. After generating the embeddings, these methods mainly focus on modeling temporal dependencies. (2) Token-wise (Liu et al., 2024a; Wang et al., 2024c): These methods treat each channel as a token and input the sequence of tokens into a Transformer model. This helps reduce irrelevant noise for forecasting. Although multi-layer Transformers have the potential to capture complex hierarchical inter-channel structures, they are extremely time-consuming and do not scale well to high-dimensional time series. Moreover, the lack of an explicit mechanism for learning hierarchical channel structure may limit their effectiveness. (3) Cluster-wise (Chen et al., 2024; Qiu et al., 2024b): These approaches divide the multivariate time series into disjoint channel clusters using clustering algorithms. However, using a single-layer clustering step may overlook the hierarchical organization present in high-dimensional time series, which can lead to suboptimal performance.